# Unsourced Random Access: A Recent Paradigm for Massive Connectivity

### v1

@@ II-C Uncoordinated Multiple Access
Figure 3: Throughput comparison of pure and slotted ALOHA.
Although slotted ALOHA doubles the throughput of pure ALOHA, it is still less than 40 %, which is limited. With the motivation to increase the throughput, several variants of ALOHA are proposed in the recent literature. For instance, in contention resolution diversity slotted ALOHA (CRDSA) [27], two copies of the packets are transmitted in two different slots, and in the case that a packet is successfully recovered in one slot, the effect of its replica in the other slot is subtracted. In this way, the maximum throughput can be increased to 52%. This scheme is improved by irregular repetition slotted ALOHA (IRSA) [26], where the packets are repeated a different number of times rather than twice. With the optimization of the repetition distributions by density evolution techniques, IRSA achieves $T_{e}\simeq 0.8$ in practical implementations. An extension of IRSA exploiting forward error correction is called coded slotted ALOHA (CSA) [25], which improves the efficiency even further, approaching a throughput of almost 1 when sufficiently low coding rates are employed. IRSA and CSA are further studied in [29, 30, 31, 32]. Some works also investigate the downlink transmission in an uncoordinated random access setup. For instance, in [53], the downlink of a massive random access setup is considered where the users recover their intended messages from a common message sent by the BS. It is assumed that the BS knows which subset of users are active, however, it is not a valid assumption in an unsourced random access setup.
Figure 4: An illustration of the sporadic user activity.
@@ III URA: A Recent Paradigm
In some massive MTC (or M2M) applications, the total number of devices can be in the order of millions, and they are only active very sporadically as illustrated in Fig. 4. This makes achieving any level of coordination infeasible. As a result, none of the multiple access techniques summarized in the previous section is suitable for enabling the communication of these devices. Specifically, the complexity of the grant-free multiple access is prohibitive due to the codebook size when each potential user is assigned a unique preamble. On the other hand, ALOHA leads to excessive collisions, significantly degrading communication efficiency.
@@
Considering a URA model with $M_{r}$ receive antennas, in which $K_{a}$ out of $K_{T}$ users are active at a given frame of length $n$, the received signal matrix in the absence of synchronization errors is written as [38]
@@
with
@@
$\left\{\begin{matrix}\mathbf{h}_{i}=1,\ \ M_{r}=1\ \ \ \ \ \ \ \ \ \ \ \ \ &\text{GMAC\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ }\\ \mathbf{h}_{i}\sim\mathcal{CN}(0,1),\ \ M_{r}=1&\text{Fading channel \ \ \ \ \ \ \ \ \ }\\ \mathbf{h}_{i}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M}),\ \ M_{r}>1&\text{MIMO fading channel}\\ \end{matrix}\right.$ (2)
@@
where $\mathbf{x}_{i}(\mathbf{u}_{i})$ is the length $n$ signal corresponding to the message bit sequence $\mathbf{u}_{i}\in\{0,1\}^{B}$ of the $i$-th user, $B$ denotes the number of information bits, and elements of the additive Gaussian noise ${\mathbf{Z}}\in\mathbb{C}^{M_{r}\times n}$ are drawn from $\mathcal{CN}({0},\sigma_{z}^{2})$. Each user selects its message index uniformly from the set $\{1,2,...,2^{B}\}$. The objective in URA is to determine $\mathbf{u}_{i}\ \forall i=1,2,...,K_{a}$ by analyzing the received signal $\mathbf{Y}$.
The existing literature on URA is reviewed in the following sections. However, to address specific constraints within the URA paradigm, we first summarize certain challenges that need to be overcome; namely, the development of bespoke multiuser coding schemes, mitigation of the heavy interference due to the random access nature and the high number of active users, balancing the computational complexity with system performance, and the reduction of preamble/pilot collisions.
III-A Development of Multiuser Coding Schemes
For encoding different users’ messages given the signal model in (1), the most immediate solution is for the active users to map their bit sequences to the rows of a $2^{B}\times n$ codebook in order to select their transmitted signals. The receiver can separate the signals of different users and detect each using detectors such as maximum likelihood (ML) [36]. Nevertheless, it is crucial to note that this coding scheme is only practical for lower values of message length $B$ since its computational complexity is exponential in $B$.
More computationally efficient approaches based on coded compressed sensing and tensor-based techniques (which will be discussed in the following sections) are adopted in [55, 54, 56, 58, 59, 60, 62, 63, 57, 61]. They involve dividing each user’s bit sequence into distinct segments, allowing each segment to be mapped to a codebook with reduced dimensions. This reduction in dimensionality helps lower the computational load at the decoder, albeit at the cost of reduced performance.
In some other URA coding schemes, each user can encode its entire message sequence (without partitioning it) with strong single-user forward error correction (FEC) codes such as polar or low-density parity-check (LDPC) codes. At the receiver side, estimation techniques such as minimum mean square error (MMSE) are employed to separate the received signals of different users, followed by single-user channel decoding [64, 65, 66, 67, 71, 68, 69, 70, 72, 73, 74, 75, 76, 37, 77, 78, 39, 79, 80, 81, 82, 38]. These algorithms are considered efficient since they manage to strike a balance, avoiding both the high computational complexity associated with ML-based coding schemes and the low accuracy observed in the schemes that segment bit sequences.
III-B Dealing with Heavy Interference
@@
Interference stands out as the primary challenge in multiple access channels. Failure to incorporate strategies for reducing or canceling it can lead to significant performance degradation. To demonstrate the detrimental impact of interference, we study the performance of the treating interference as noise (TIN) scheme in a MIMO fading setup, where single-user decoding is carried out by taking interference from other users as noise, without employing any special technique to mitigate it. Approximating $\frac{1}{M_{r}}\|\mathbf{h}_{i}\|^{2}\approx 1\ \forall i$ in (1), the signal-to-interference-plus-noise ratio (SINR) of the $i$-th user, $\alpha_{i}$, in the TIN strategy can be written as
$\alpha_{i}\approx\dfrac{P}{\sigma_{z}^{2}+{(K_{a}-1)}P},$ (3)
where $P$ is the transmitted power of each user. For large enough active user load $K_{a}$, the SINR can be written as
@@
This equation illustrates that when the number of users surpasses a specific threshold, the SINR becomes inversely proportional with $K_{a}$, and boosting the transmit power does not result in an enhanced SINR. Consequently, a significant rise in $K_{a}$ (equivalent to an increase in interference) adversely impacts the system’s performance. Furthermore, one can observe the huge performance gap between TIN and the SIC-added version of it (TIN-SIC) via the results in [83], where the TIN scheme is compared with TIN-SIC as a strategy designed to mitigate interference.
There are some common techniques in the URA literature to mitigate the interference:
@@
    Slotted transmission: Dividing the time frame into multiple slots serves to diminish interference by reducing the number of users in each slot [83, 84, 82, 38]. Additionally, slotted transmission results in a shorter duration for the transmitted signal and a reduction in the number of interfering users per slot. Consequently, the overall computational complexity of the URA system is decreased since an increase in these two parameters superlinearly increases the computational workload of the multiuser decoder [39, 38].
@@
    Sending some preambles/pilots: In some URA schemes, each user randomly selects a preamble/pilot, and transmits it along with its main signal [39, 37, 76, 66, 67, 81, 82, 38]. Upon detecting these transmitted preambles/pilots at the receiver, there is an opportunity to signify the influence of a desired user while mitigating the impact of other interfering ones. For instance, in certain MIMO fading URA schemes, each user selects a pilot and appends it to their codeword before transmission. At the receiver end, the channel coefficients for different users can be estimated using the detected pilots. With these estimated channel vectors for various users, interference can be significantly reduced by employing estimators such as MMSE and maximum ratio combining (MRC) [37, 76, 82, 38, 39].
@@
    SIC: Successive interference cancellation (SIC) is a useful technique for interference mitigation in wireless multiuser scenarios. The idea behind it is to re-encode the successfully detected messages and remove their effects from the received signal to decrease the interference [85, 86, 81, 82, 38]. Note that in fading channels with unknown CSI, the overall performance can be improved with a simple additional step where the corresponding channel coefficients for the user of a detected message are re-estimated before proceeding with the subtraction process [82, 38].
  4. 4.
    Power diversity: An alternative approach for user separation involves randomly distributing users into different groups with varying power levels and adopting an appropriate technique to ensure that the decoding process for different groups is independent of each other [81, 38]. This approach serves to reduce interference within each group. As previously discussed in the context of slotted transmissions, dividing users into distinct groups also leads to a decrease in the computational complexity of the URA system.
III-C Addressing Computational Complexity
To increase the number of active users supported by a URA system, larger values of parameters such as frame length and the number of antenna elements in MIMO systems, are necessary [36, 58, 87]. Additionally, the computational complexity of a URA system escalates with these parameters and the number of active users, resulting in a superlinear dependence of computational load on the active user count in most URA solutions [82, 39]. Consequently, a URA system with a substantial number of active users is anticipated to exhibit a high computational complexity. To mitigate this complexity, various strategies are employed, such as slotted transmissions, and randomly dividing users into different independent groups [82, 39].
III-D Reducing Pilot Collisions
@@
Many existing fading URA solutions employ preambles/pilots to estimate the channel coefficients used in the decoding phase. However, it is impossible to assign dedicated preambles/pilots due to the massive number of potential users. Therefore, as a common approach, the users randomly and independently pick a preamble/pilot from a common codebook based on a part of their message bits (pilot bits). Hence, if the pilot bits of multiple users are the same (i.e., pilot contamination), a collision occurs and the system struggles to decode the transmitted packets. Therefore, it is crucial to perform a preliminary analysis to optimize the design parameters to keep the collision probability lower than a desired threshold [38] or develop techniques for its mitigation [82].
@@ IV URA over GMAC
IV-A Achievability Bounds
In [36], Polyanskiy introduces the URA paradigm and obtains an achievability bound on its performance on the following GMAC model:
@@ IV-A Achievability Bounds
where $K_{a}$ is the number of active users (i.e., the active user load), $\mathbf{x}_{i}$ is the transmitted signal of the $i$-th user, $\mathbf{y}$ is the received signal and $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$ is the AWGN noise vector with independent and identically distributed (i.i.d.) elements. Assuming $\norm{\mathbf{x}_{i}}^{2}\leq nP$ is satisfied, the energy-per-bit becomes (actually, it is upper bounded by)
$\frac{E_{b}}{N_{0}}=\frac{nP}{2\log_{2}M},$ (6)
where $M$ is the cardinality of the set of the possible messages $[M]$ and $n$ is the length of the transmission frame. In [36], an $(M,n,\epsilon)$ random access code is defined as follows
@@
An $(M,n,\epsilon)$ random access code for the $K_{a}$-user channel $P_{Y\mid X^{K_{a}}}$ is a pair of (possibly randomized) maps, the encoder $f:[M]\rightarrow\mathcal{X}^{n}$ and the decoder $g:\mathcal{Y}^{n}\rightarrow\binom{[M]}{K_{a}}$, satisfying
@@
The achievability bound obtained in [36], which is given below, has been used as a benchmark for the energy efficiency of URA over a real-valued AWGN channel.
Theorem 1 (Random coding achievability [36]).
@@
$q_{t}=\inf_{\gamma}\mathbb{P}[I_{t}\leq\gamma]+exp(n(tR_{1}+R_{2})-\gamma).$ (9)
@@
where $C_{t}=\frac{1}{2}\log(1+P^{\prime}t)$. $\blacksquare$
The idea behind the proof of this theorem is as follows: Let $S=\{W_{1},\dots,W_{K_{a}}\}$ be a random $K_{a}$-subset of $[M]$. We can bound the error event $F_{t}\triangleq\{\lvert S\setminus\hat{S}\rvert=t\}$ with two different methods. The first one is a Gallager-type bound, defining the error events using the generic subsets of size $t$ in $[K_{a}]$ and $[M]\backslash[K_{a}]$, respectively, using the Chernoff bound and applying Gallager’s $\rho$-trick twice. The second one utilizes the information density in (11), considering the event $\tilde{F}=\{I_{t}<\gamma\}$ with the fixed $\gamma$ and $I_{t}$ is given in (10), and taking union over $S_{0}$. The details can be found in [36].
The random coding bound is an information-theoretic performance limit, and it is derived without any complexity constraints as it assumes that all active users are jointly decoded, which has a prohibitive complexity. So, practical coding schemes for URA with reasonable complexity are needed to achieve the results predicted by the bound.
It is assumed in slotted ALOHA that the decoding works if there is no collision and the single-user decoding is successful as the collided packets are discarded in slotted ALOHA protocol. Moreover, as shown in [36], TIN coding satisfies the following approximation
@@
where $C_{TIN}(P)=\frac{1}{2}\log\left(1+\frac{P}{1+(K_{a}-1)P}\right)$.
Figure 5: Required $E_{b}/N_{0}$ versus the number of active users.
@@
In [36], the performances of these approaches are evaluated assuming that there are $K_{a}$ active users transmitting $k=100$ bits of information through a transmission frame of length $n=30000$, with a target PUPE of $\epsilon=0.1$. Note that the same set of parameters has been adopted by other GMAC URA solutions for performance analysis, for obvious reasons. However, a more comprehensive analysis of the existing competing solutions, especially those nominated for inclusion in future protocols, is needed that covers different sets of parameters representing different future URA applications.
@@
The required $E_{b}/N_{0}$ versus number of active users for these schemes are plotted in Fig. 5 along with the random coding bound derived in [36] and a converse bound for $K_{a}$-GMAC from [88]. The results in Fig. 5 show that the conventional random access techniques (i.e., slotted ALOHA and TIN) are energy inefficient for the URA scenario as there is a huge performance gap between these and the random coding bound, and the development of low-complexity coding schemes is needed. The performance of the grant-based TDMA scheme where the frame is divided into sub-frames of length $\frac{n}{K_{a}}$ and each active user transmits in its dedicated sub-frame referred to as Orthogonal M.A. (multiple access) is also plotted in the figure (though being infeasible for URA). Note that in all the curves, it is assumed that the single-user codes satisfying the normal approximation in [89] are employed.
@@
In [36], the random coding bound is developed under the assumption that the receiver knows the number of active users $K_{a}$. The authors of [90] extend the random coding bound in [36] to the case that the number of active users is random and unknown at the receiver. Furthermore, an achievability bound for the binary-input signaling for URA over GMAC is derived in [91]. A random coding bound for the scenario that the users send a common alarm message in addition to standard messages is developed in [92], where the alarm message needs to be decoded with a lower error rate than the standard ones. Moreover, an extension to the case of asynchronous GMAC is provided in [93], where the delay in transmission can be up to a small portion of the frame. However, the provided achievability bound is a loose bound for the asynchronous case as the authors have employed a fixed time-frame decoder in which all the packets that have some overlaps with the end of the frame are deemed undecodable, leading to huge performance degradation.
IV-B Existing Solutions for GMAC URA

### v2

@@ II-B1 ALOHA & Variants
With the motivation to increase the throughput, several variants of SA are proposed in the recent literature, where the main idea is to transmit multiple copies of a packet in different slots. If it is successfully recovered in one slot, the effect of its replicas in the other slots is removed. Based on this idea, two copies of the same packet is transmitted in contention resolution diversity slotted ALOHA (CRDSA) [27], and the packets are repeated a different number of times rather than twice in irregular repetition slotted ALOHA (IRSA) proposed in [28]. An extension of IRSA exploiting forward error correction that is called coded slotted ALOHA (CSA) [29] approaches a throughput of almost one when sufficiently low coding rates are employed. IRSA and CSA are further studied in [30, 31, 32, 33, 34]. Moreover, the resolution of partial collisions in random access setups is discussed in [35].
II-B2 Grant-Free Random Access
Grant-free RA is another RA candidate to enable communication of a large number of users where the users transmit their data by employing the shared system resources without any prior grant from the BS, which reduces the latency and signaling overhead. Nevertheless, the users can have unique signatures/preambles or they can utilize different codebooks, which means that there is user identity during the data transmission phase, and the receiver can utilize this information for activity detection or user separation [36]. This is called sourced (grant-free) random access, which one of solutions based on it, grant-free NOMA, is the prominent solution for future communication systems due to its low latency, improved spectral efficiency, and potential to support massive connectivity. Note that the main difference between grant-based and grant-free NOMA is that, the BS does not have information on the set of active users in the latter, which brings significant challenges requiring different solutions such as blind activity detection, channel estimation, and data recovery, and synchronization with minimal overhead [37].
Among the two main categories of NOMA, the power-domain NOMA is hard to extend to the grant-free scenario as the successful recovery of the transmitted packets at the receiver heavily depends on the power difference of the users, which may not be maintained due to the grant-free nature of the transmissions. On the other hand, signature-based schemes in code-domain NOMA can be adapted to the grant-free case with proper activity detection before the data detection or joint activity detection and data recovery [38]. For instance, a grant-free version of SCMA is proposed in [39]. Furthermore, the inherent sparsity of the user activity in grant-free MA is utilized in [40, 41, 42, 43] through compressive sensing algorithms for multiuser detection.
@@ III URA: A Recent Paradigm
In some next-generation communications applications, such as massive MTC and massive IoT, the total number of devices can be in the order of millions, and they are only active very sporadically, as illustrated in Fig. 3. This makes achieving any level of coordination infeasible. As a result, none of the MA techniques summarized in the previous section is suitable for enabling the communication of these devices. Specifically, the complexity of the grant-free MA is prohibitive due to the codebook size when each potential user is assigned a unique preamble. On the other hand, ALOHA leads to excessive collisions, significantly degrading communication efficiency.
@@
III-A General System Model
The URA system model in a synchronous setup as follows. Considering a URA system with an $M_{r}$-antenna receiver and single-antenna users, in which $K_{a}$ out of $K_{T}$ users are active at a given time frame of length $n$, the received signal, $\mathbf{Y}\in\mathbb{C}^{M_{r}\times n}$, in the absence of synchronization errors can be written in the following general form [13]
@@ III-A General System Model
with the following common models in the URA literature
@@
$\begin{cases}M_{r}=1,\mathbf{h}_{i}=1,&\text{GMAC}\\ M_{r}=1,\mathbf{h}_{i}\sim\mathcal{CN}(0,1),&\text{single-antenna fading channel}\\ M_{r}>1,\mathbf{h}_{i}\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M_{r}}),&\text{MIMO fading channel}\\ \end{cases}$ (2)
@@
where $\mathbf{x}_{i}(\mathbf{u}_{i})$ is the length-$n$ transmitted signal of the $i$-th user corresponding to its message bit sequence $\mathbf{u}_{i}\in\{0,1\}^{B}$ of length $B$, and ${\mathbf{Z}}\in\mathbb{C}^{M_{r}\times n}$ is the circularly-symmetric additive white Gaussian noise (AWGN) whose elements are drawn from $\mathcal{CN}({0},\sigma_{z}^{2})$, with $\sigma_{z}^{2}$ being the power of noise. In the literature, it is assumed that each user selects its message index uniformly from the set $\{1,2,...,2^{B}\}$. The objective in URA is to determine a list of transmitted messages, $\mathbf{u}_{i}^{\prime}s,\ \forall i=1,2,...,K_{a}$, by processing the received signal $\mathbf{Y}$.
Figure 3: An illustration of a system where a large number of potential devices (users) try to communicate with an access point (AP), with only a small subset of them being active at each time frame.
In the URA literature, the common setup involves users having messages of fixed length $B$ bits from a set of the possible messages $[M]$ with cardinality $M$, and the objective is to minimize the required energy-per-bit for a target per-user probability of error (PUPE), as the performance metric. The energy-per-bit can be written as
$\dfrac{E_{b}}{N_{0}}=\dfrac{nP}{c\sigma_{z}^{2}B},$ (3)
where $P$ denotes the average power of each user per channel use, i.e., $\norm{\mathbf{x}_{i}}^{2}\leq nP$, and we have $c=2$ for real-valued and $c=1$ for complex channels. Also, PUPE is defined in terms of the probability of false-alarm, $p_{fa}$, and the probability of misdetection, $p_{md}$, as
$P_{e}=p_{fa}+p_{md},$ (4)
where $p_{md}=\dfrac{1}{K_{a}}\sum_{i\in\mathcal{K}_{a}}{\mathrm{Pr}(\mathbf{u}_{i}\notin\mathcal{L}_{d})}$, where $\mathcal{K}_{a}$ and $\mathcal{L}_{d}$ are the set of active user indices and the list of decoded messages, respectively, and $p_{fa}=\mathbb{E}\left\{\dfrac{n_{fa}}{|\mathcal{L}_{d}|}\right\}$, with $n_{fa}$ being the number of decoded messages that were indeed not sent.
III-B Challenges
In the following sections, we review the existing URA solutions in the literature. However, in order to have a better understanding of the URA setup and the reasoning behind the provided solutions in the literature, we first need to know the main challenges faced in a URA setup, which will be hinted at in the rest of this section.
@@ III-B Challenges
Interference: Interference stands out as the primary challenge in MA channels, especially random access due to less coordination. It has an even more detrimental effect on the URA setup since there is no coordination between the receiver and the users. Failure to incorporate strategies for reducing or canceling it can lead to significant performance degradation. To demonstrate the detrimental impact of interference, we analyze the treating interference as noise (TIN) scheme, where single-user decoding is carried out by taking interference from other users as noise, without employing any special technique to mitigate it. The signal-to-interference-plus-noise ratio (SINR) of the $i$-th user, $\alpha_{i}$, in the TIN strategy for a GMAC scenario can be written as
$\alpha_{i}=\dfrac{P}{\sigma_{z}^{2}+{(K_{a}-1)}P},$ (5)
[…]
@@
This equation illustrates that when the number of users surpasses a specific threshold (more specifically, when it is much greater than the inverse of transmit signal-to-noise ratio (SNR), i.e., $K\gg\frac{\sigma^{2}_{z}}{P}$, which is not a very large value in many envisioned URA applications), the SINR becomes inversely proportional with $K_{a}$. So, boosting the transmit power does not result in an enhanced SINR or, equivalently, the system performance. Consequently, in systems such as URA, the relatively large number of active users significantly rise in $K_{a}$ (equivalent to an increase in interference) adversely impacts the system performance.
There are two main techniques in the URA literature to alleviate the effects of interference, with the most common one being SIC. SIC is a useful technique for interference mitigation in wireless multiuser scenarios, which removes the effects of successfully detected messages from the received signal to decrease the interference before another round of decoding. Especially in URA, SIC is an essential step in achieving low error rates due to heavy multiuser interference. The other technique for interference management involves randomly separating users into different groups, e.g., separation in time by slotting the time frame, in the power domain by employing different transmit power levels, or in the code domain by multiplying or concatenating a known sequence to the data segment.
@@
    Time-domain separation: Dividing the time frame into multiple slots and letting each user randomly select a slot to transmit its data can lead to enhanced system performance. In a slotted structure, the trade-off is between the effects of decreased interference due to the reduced number of interfering users in each slot and decreased channel coding power due to the fewer channel uses available to each user. Therefore, optimizing the number of slots as a function of system parameters such as active user load plays an important role in benefiting from the slotting.
@@
    Code-domain separation: In some URA schemes, each user randomly selects a preamble/pilot/pattern and uses it to prepare its data for transmission. One way is to transmit a concatenation of a preamble or pilot and the main signal. At the receiver end, the channel coefficients for different users can be estimated using the detected pilots. Then, minimum mean square error (LMMSE) or maximum ratio combining (MRC) solutions can be employed to separate the user packets, before decoding them using single-user channel decoders. Another way is to multiply a randomly selected pattern with the main signal, which is used in random-spreading-based solutions. Due to limited available power and time resources, employing these techniques usually decreases the amount of power and time resources dedicated to the data (main signal) part. So, in order for them to be effective in interference management, the choices of type, length, and power of the pilots and pattern sequences are of great importance.
@@
    Power-domain separation: An alternative approach for user separation, especially effective in GMAC scenarios, is the so-called power diversity technique [44], where each user randomly picks a transmit power level from a limited set with a certain probability distribution, satisfying the overall power constraint. Since the users are now separated by different power levels, the receiver, with a high probability, decodes the users with the highest power level first, and after each SIC, it moves to decode the lower power levels. The appropriate design of the power diversity parameters, namely, the number of power levels, their values, and their probability distribution, is very important in achieving the power diversity gain.
Collisions: In many classic random access setups, such as ALOHA, collision is used interchangeably with interference, i.e., when user packs are fully or partially received at the same time at the receiver. However, in the URA literature, collision by default means the collision of either user messages or user pilots, which are when two or more users have exactly the same messages or pilot sequences, respectively.
  1. 1.
    Message collision: In the case of user message collision, with the help of a randomized encoding process independent of user messages, the receiver can recover collided user messages to some extent. Note that in a conventional URA setup, since the receiver is only interested in the list of transmitted messages, being able to recover at least one of the collided messages does not increase the misdetection rate. However, this matter becomes more important in some URA systems where the receiver is also interested in the frequency of user messages.
  2. 2.
@@
    Pilot collision: Many existing fading URA solutions employ preambles/pilots to estimate the channel coefficients used in the decoding phase. However, it is impossible to assign dedicated preambles/pilots due to the massive number of potential users. Therefore, as a common approach, the users independently pick a preamble/pilot from a common codebook, either randomly or based on a part of their message bits. Hence, if the pilot bits of multiple users are the same (called pilot contamination in massive MIMO literature [45]), a collision occurs, and the system struggles to decode the transmitted packets. Therefore, it is crucial to perform a preliminary analysis to optimize the design parameters to keep the pilot collision probability lower than a desired threshold or develop techniques for its mitigation.
Computational Complexity: In order to increase the number of active users supported by a URA system, larger values of parameters such as time frame length and the number of antenna elements (in MIMO systems) are necessary. However, the computational complexity of URA solutions escalates superlinearly with these parameters as well as the number of active users Consequently, solutions for a URA system with a high dimension, temporal (time frame length), spatial (number of receiver antennas), and/or the number of active users, are anticipated to exhibit a high computational complexity. To mitigate this complexity, various strategies are employed in the literature, such as slotted transmissions and coded compressed sensing (CCS). The idea of CCS is to divide the bit sequence into smaller segments, map each segment to a column of a common codebook, and stitch them using a tree code instead of mapping the entire bit sequence to the columns of a $2^{B}\times n$ codebook, which would lead to a prohibitive receiver complexity. This reduction in dimensionality helps lower the computational load at the decoder, albeit at the cost of reduced performance.
In this section, we summarized the main challenges in a URA setup along with a high-level discussion on the proposed solutions in the literature. In the two subsequent sections, we take a more detailed look at these solutions, with Section IV dedicated to GMAC and Section V to single-antenna (V.A) and MIMO (V.B) fading setups. We take a systematic approach in presenting the URA solutions in these sections: different contributions are grouped based on their solution approach. In each group, the underlying approach is explained with an emphasis on the main work accompanied by a block diagram of its encoder and decoder. The studies are presented in chronological order. To provide an overall view, at the end of each main subsection, a summary of the primary contributions is presented, and a comparative discussion of their performance is provided, accompanied by some numerical results.
@@ IV URA over GMAC
In this section, we present an overview of GMAC URA. To this end, we first present an achievability bound on the performance of the GMAC URA, as presented in Polyanskiy’s original paper [11], and then, review the existing GMAC URA solutions to achieve it.
IV-A Performance Bounds
In [11], Polyanskiy introduces the URA paradigm and obtains an achievability bound on its performance on the following GMAC model with single-antenna users and receiver:
@@ IV-A Performance Bounds
where $K_{a}$ is the number of active users (i.e., the active user load), $\mathbf{x}_{i}$ is the transmitted signal of the $i$-th user, $\mathbf{y}$ is the received signal and $\mathbf{z}$ is the circularly-symmetric AWGN noise vector with independent and identically distributed (i.i.d.) elements, i.e., $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n})$.
In [11], an $(M,n,\epsilon)$ RA code, where $\epsilon$ is the target PUPE, is defined as follows
@@
An $(M,n,\epsilon)$ RA code for the $K_{a}$-user channel $P_{Y\mid X^{K_{a}}}$ is a pair of (possibly randomized) maps, the encoder $f:[M]\rightarrow\mathcal{X}^{n}$ and the decoder $g:\mathcal{Y}^{n}\rightarrow\binom{[M]}{K_{a}}$, satisfying
@@
The following achievability bound has been used as a benchmark for the energy efficiency of URA over a real-valued AWGN channel.
Theorem 1 (Random coding achievability bound [11]).
@@
$q_{t}=\inf_{\gamma}\mathbb{P}[I_{t}\leq\gamma]+exp(n(tR_{1}+R_{2})-\gamma),$ (10)
@@
where $C_{t}=\frac{1}{2}\log(1+P^{\prime}t)$.
The idea behind the proof of this theorem is as follows: Let $S=\{W_{1},\dots,W_{K_{a}}\}$ be a random $K_{a}$-subset of $[M]$. We can bound the error event $F_{t}\triangleq\{\lvert S\setminus\hat{S}\rvert=t\}$ with two different methods. The first one is a Gallager-type bound, defining the error events using the generic subsets of size $t$ in $[K_{a}]$ and $[M]\backslash[K_{a}]$, respectively, using the Chernoff bound and applying Gallager’s $\rho$-trick twice. The second one utilizes the information density in (12), considering the event $\tilde{F}=\{I_{t}<\gamma\}$ with the fixed $\gamma$ and $I_{t}$ is given in (11), and taking union over $S_{0}$.
The random coding bound is an information-theoretic performance limit, and it is derived without any complexity constraints as it assumes that all active users are jointly decoded, which has a prohibitive complexity. So, practical coding schemes for URA with reasonable complexity are needed to achieve the results predicted by the bound. In order to illustrate the ineffectiveness of common RA solutions in a URA setup, we first compare their performance with Polyanskiy’s URA achievability result.
In slotted ALOHA, a user packet is successfully recovered if there is no collision and the single-user decoding is successful, as the collided packets are discarded. Moreover, using the normal approximation in [46], the finite blocklength capacity of TIN can be approximated as
@@
where $C_{TIN}(P)=\frac{1}{2}\log\left(1+\frac{P}{1+(K_{a}-1)P}\right)$, $Q^{-1}(\cdot)$ denotes the inverse Q-function.
@@
In [11], the performances of these approaches are evaluated assuming that there are $K_{a}$ active users transmitting $B=100$ bits of information through a transmission time frame of length $n=30000$, with a target PUPE of $\epsilon=0.1$. Note that the same set of parameters has been adopted by subsequent GMAC URA works for performance comparison, for obvious reasons. However, a more comprehensive analysis of the existing competing solutions, especially those nominated for inclusion in future protocols, is needed that covers different sets of parameters representing different future URA applications.
@@
The required $E_{b}/N_{0}$ versus number of active users for these schemes are plotted in Fig. 4 along with the random coding bound derived in [11] and a converse bound for $K_{a}$-GMAC from [47]. The results in Fig. 4 show that the conventional RA techniques (i.e., slotted ALOHA and TIN) are energy inefficient in a URA setup as there is a huge performance gap between these and the random coding bound, and the development of low-complexity coding schemes is needed. The performance of the grant-based TDMA scheme is also plotted in the figure (referred to as Orthogonal M.A) as a benchmark, in which the time frame is divided into sub-frames of length $\frac{n}{K_{a}}$ and each active user transmits in its dedicated sub-frame, with its required $E_{b}/N_{0}$ being calculated using the normal approximation in [46]. However, note that the TDMA scheme is not feasible in a URA setup, mainly due to the lack of coordination. Putting the coordination issue aside, the system works in a regime of $n<K_{a}B$ in some URA applications, which renders orthogonal solutions infeasible.
Figure 4: Performance comparison of conventional random access techniques with Polyanskiy’s URA (random coding) achievability bound.
@@
The random coding bound in [11] is developed under the assumption that the receiver knows the number of active users $K_{a}$. The authors of [48] extend this bound to the case that the number of active users is random and unknown at the receiver. Furthermore, an achievability bound for the binary-input signaling for URA over GMAC is derived in [49]. A random coding bound for the scenario that the users send a common alarm message in addition to standard messages is developed in [50], where the alarm message needs to be decoded with a lower error rate than the standard ones. This result is extended to the case of asynchronous GMAC in [51], where the delay in transmission is assumed to be up to a small portion of the time frame. However, the provided achievability bound is a loose bound for the asynchronous case as the authors have employed a fixed time frame decoder in which all the packets that have some overlaps with the end of the time frame are deemed undecodable, leading to huge performance degradation, which is not valid in a fully asynchronous setup, where the time horizon is unbounded. Furthermore, the problem of joint delay and activity detection in the asynchronous setup is investigated in [52] using methods of tropical geometry, which is the study of the geometric properties of polynomials when addition is replaced by minimization and multiplication by ordinary addition. Moreover, the conditions for path delays for zero-error detection are identified.
IV-B URA Solutions for GMAC